# Understanding Federated Learning from IID to Non-IID dataset: An Experimental Study

### v2

@@ 2 From Gradient Descent to FedAvg
To balance the trade-offs between GD and SGD, mini-batch SGD was developed. It computes gradients using a batch of size $B$ for each update, reducing SGD’s instability while improving GD’s efficiency. GD, SGD, and mini-batch SGD can be seen as special cases within a broader framework, where the batch size $B$ ranges from 1 (SGD) to the entire dataset (GD), as shown in Algorithm LABEL:alg.sgd, with their respective optimization paths illustrated in Fig. 1.
The concept of gradient averaging for training extends naturally to distributed and parallel computing, as gradient calculations are independent until aggregation. Scaling SGD process across multiple devices (e.g., GPUs or distributed systems) accelerates training and mitigates the limitations of centralized hardware. While there are various approaches to parallel SGD [3, 23, 4], in this work, we focus on aggregating gradients from each device per batch of each client, as described in Algorithm 1. In this distributed learning approach, we introduce a parameter, $K$, representing the number of devices. Gradients are computed on each device, followed by averaging locally and then globally.

### v3

@@ 2 From Gradient Descent to FedAvg
To balance the trade-offs between GD and SGD, mini-batch SGD was developed. It computes gradients using a batch of size $B$ for each update, reducing SGD’s instability while improving GD’s efficiency. GD, SGD, and mini-batch SGD can be seen as special cases within a broader framework, where the batch size $B$ ranges from 1 (SGD) to the entire dataset (GD), as shown in Algorithm 1, with their respective optimization paths illustrated in Fig. 1.
Algorithm 1 Stochastic Gradient Descent (SGD)
1:  Input: Learning rate $\eta$, Batch size $B$, Number of epochs $E$
2:  for $n=1$ to $E$ do
3:     for each random mini-batch $\mathcal{B}$ of size $B$ with data $(x_{i},y_{i})$ do
4:        Compute gradient $g=\frac{1}{B}\sum_{i\in\mathcal{B}}\nabla_{\theta}F(\theta;x_{i},y_{i})$
5:        Update parameters: $\theta\leftarrow\theta-\eta\cdot g$
The concept of gradient averaging for training extends naturally to distributed and parallel computing, as gradient calculations are independent until aggregation. Scaling the SGD process across multiple devices (e.g., GPUs or distributed systems) accelerates training and mitigates the limitations of centralized hardware. While there are various approaches to parallel SGD [3, 23, 4], in this work, we focus on aggregating gradients from each device per batch of each client, as described in Algorithm 2. In this distributed learning approach, we introduce a parameter, $K$, representing the number of devices. Gradients are computed on each device, followed by averaging locally and then globally.